OpenStudy (anonymous):

What is 5/6 times 3/10? A. 1/16 b. 1/8 c. 1/4 d. 1/2

OpenStudy (anonymous):

1/4, or 0.25

OpenStudy (anonymous):

So c, bassicaly

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Shoot :)

OpenStudy (anonymous):

What is 3/14 divided by 4/7?

OpenStudy (anonymous):

6/49

OpenStudy (anonymous):

This is fun, more?

OpenStudy (anonymous):

thanks!! yess more

OpenStudy (anonymous):

Okay, shoot :D

OpenStudy (anonymous):

wait 6/49 isn't one of the answer choices... Thers's just 3/4 ,3/8, 2 2/3 and 12/21

OpenStudy (anonymous):

Hmmm... hold on

OpenStudy (anonymous):

3/8! My bad, thought you said multiply!

OpenStudy (anonymous):

I had a feeling you thought I said that!! Thanks!! More??

OpenStudy (anonymous):

Yes! Shoot!

OpenStudy (anonymous):

13/15 plus -1/5? A.1/2 B.2/3 C.1 1/5 D.3/5

OpenStudy (anonymous):

it's 2/3 :D (b)

OpenStudy (anonymous):

thank you!! im horrible in math

OpenStudy (anonymous):

No problem cx was that it?

OpenStudy (anonymous):

What is 3 5/6 minus 1/3?

OpenStudy (anonymous):

7/2 :P

OpenStudy (anonymous):

A.3 1/3 , B.3 1/2, C.3 2/3, D.4 1/3

OpenStudy (anonymous):

Hmmm... gimme a sec

OpenStudy (anonymous):

That's weird... I never get anything but 7/2

OpenStudy (anonymous):

Wait! It's either A or D...

OpenStudy (anonymous):

Did you do 3 and 5/6 minus 1/3? or did you just do 5/6-1/3?

OpenStudy (anonymous):

I believe it's A (yes I did)

OpenStudy (anonymous):

ok thanks, that was a tricky one

OpenStudy (anonymous):

Yes it was... Anything else?

OpenStudy (anonymous):

What is the solution to the equation? 2/5x =20? A.10 B.40 .C.50 D.8

OpenStudy (anonymous):

I think it's A

OpenStudy (anonymous):

C, acctualy

OpenStudy (anonymous):

I got this one @JungHyunRan, go help someone else

OpenStudy (anonymous):

\(\Large \frac{2}{5x}=20 \Rightarrow x=\frac{2}{20\times5}=\frac{1}{50}\)

OpenStudy (anonymous):

(Didn't mean too sound rude cx)

OpenStudy (anonymous):

Basicaly, yes. It is C

OpenStudy (anonymous):

Ohh ok hmm , that was funny what you said to JungHyun

OpenStudy (anonymous):

Yeah, my bad, sounded kind of rude... But I'm on a role with your question xD

OpenStudy (anonymous):

It's \(\Large \frac{2}{5x} \) or \(\Large \frac{2x}{5}\)??

OpenStudy (anonymous):

*questions

OpenStudy (anonymous):

What is the least common denominator of 3/20 and 8/15? A.5 B.60 C.20 D.30

OpenStudy (anonymous):

\(\Large 60\)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Ok I thought so thanks

OpenStudy (anonymous):

Sorry for the Delay

OpenStudy (anonymous):

What is 3 1/4 times -2 3/5?

OpenStudy (anonymous):

Hey Sailor moon, go help cassious

OpenStudy (anonymous):

And btw that's 3 and 1/4 not 31/4 and it -2 and 3/5

OpenStudy (anonymous):

−91/20

OpenStudy (anonymous):

A.-8 9/20, B.-6 3/20, C.6 3/20, D.8 9/20

OpenStudy (anonymous):

\(\huge 3\frac{1}{4} \times -2\frac{3}{5}=\frac{13}{4} \times \frac{-13}{5}=\frac{-169}{20}\)

OpenStudy (anonymous):

D..? maybe

OpenStudy (anonymous):

the correct answer is : \(\Large -8\frac{9}{20}\)

OpenStudy (anonymous):

Do I put that in simpliest form? What about B?

OpenStudy (anonymous):

A? that doesn't sound right... but okay

OpenStudy (anonymous):

B isn't A simplified Jada

OpenStudy (anonymous):

umm I don't understand. What do you mean B isn't A siplified?

OpenStudy (anonymous):

What do you think it is @CA_Vegasvinny ?

OpenStudy (anonymous):

I think it's D... but A does sound partially right

OpenStudy (anonymous):

hmm interesting

OpenStudy (anonymous):

I'll tell you if i got it right or nah

OpenStudy (anonymous):

I think you should go with A, Jung is smarter then me